# Securing the Skies: A Comprehensive Survey on Anti-UAV Methods, Benchmarking, and Future Directions

### v2

@@ Abstract
Dataset | Task | Modality | Seq. | Frame | Annotation | Resolution | Rate (Hz) | T-Type | S-Type | Venue (Year) |
FL-Drones [103] | D, T | RGB | 14 | 39K | Bbox | 640480 / 752480 | – | 1 | 2 | TPAMI (2016) |
NPS-Drones [68] | D, T | RGB | 50 | 70K | Bbox | 19201280 / 1280760 | – | 1 | – | IROS (2016) |
USC Drone [24] | D, T | RGB | 60 | 20K | Bbox | 19201080 | 15 | 1 | – | APSIPA ASC (2017) |
DroneRF [8] | C, D | RF | 454 | 10M | Class | 14-bit (RF) | Less than 1K (RF) | 3 | – | Data in Brief (2019) |
DroneAudio [2] | C, D | A | – | 1332 clips | Class | – | 16K (A) | 2 | – | IWCMC (2019) |
MAV-VID [100] | D, T | RGB | – | 40K | Bbox | – | – | 1 | – | IEEE Access (2020) |
Hui [55] | D, T | IR | 22 | 16K | Centroid | 256256 | 100 | 1 | 22 | CSD (2020) |
Acoustic-UAV [21] | C, D | A | – | 91K sec | Class | – | 22.05K (A) | 9 | 6 | Drones (2021) |
mDrone [165] | C, D | RA | 100 | 60K | 3D coords | – | 10 (RA) | 1 | – | ICRA (2021) |
Det-Fly [166] | D | RGB | – | 13K | Bbox | 38402160 | 5 | 1 | 4 | RA-L (2021) |
MOT-Fly [28] | C, D | RGB | 16 | 11K | Bbox | 19201080 | – | 3 | – | ICUS (2021) |
Halmstad Drone [122] | D, T | RGB, IR, A | 650 | 203K | Bbox | 640512 (RGB), 320256 (IR) | 60 (IR), 30 (RGB) | 4 | 3 | ICPR (2021) |
Drone-vs-Bird [29] | C, D | RGB | 77 | 105K | Bbox | 720576 – 38402160 | – | 8 | – | AVSS (2021) |
Anti-UAV [58] | D, T | RGB, IR | 318 | 297K | Bbox | 19201080 (RGB), 640512 (IR) | 25 | – | 7 | TMM (2021) |
UAVSwarm [133] | D, T | RGB | 72 | 12K | Bbox | 446276 – 19191079 | – | 19 | 13 | Remote Sens. (2022) |
Synthetic Drone [140] | C | RGB | – | 2K | Class | 256256 | – | 4 | 10 | CORD (2022) |
DUT Anti-UAV [163] | D, T | RGB | 20 | 35K | Bbox | 160240 – 37445616 | – | 1 | 7 | T-ITS (2022) |
AOT [10] | D, T | IR | 4943 | 5.9M+ | Bbox | 24482048 | 10 | – | – | AWS (2023) |
Anti-UAV600 [171] | D, T | IR | 600 | 723K | Bbox | 640512 | 25 | 1 | – | Preprint (2023) |
Anti-UAV410 [51] | D, T | IR | 410 | 438K | Bbox | 640512 | 25 | 1 | 7 | TPAMI (2024) |
MMAUD [156] | C, D, T | RGB, A, RA, L | 50 | 45K | Trajectory | 2560960 (RGB) | 30 (RGB), 41.8 (A), 15 (RA), 10 (L) | 6 | – | ICRA (2024) |
RGBT-Tiny [153] | D, T | RGB, IR | 115 | 93K | Bbox | 640512 | 15 | 7 | 8 | TPAMI (2025) |
@@ 1 Introduction
Al-Sa’d et al. (2019) [3] | RF | 3 DNNs (RF fingerprint) | DroneRF [8] | Introduced DroneRF (open-source, multiple flight modes). Developed three DNN classifiers for UAV presence, type, and flight mode. Achieved 99.7% (2-class), 84.5% (4-class), 46.8% (10-class). Forms an early baseline on DroneRF. |
Al-Emadi et al. (2020) [1] | RF | CNN | DroneRF [8] | Achieved 99.8% (2-class), 85.8% (4-class), and 59.2% (10-class). Comparable to Al-Sa’d et al. [3] on 2-class tasks, showing CNN’s viability for RF-based identification. |
Medaiyese et al. (2021) [80] | RF | XGBoost pipeline | DroneRF [8] | Used 227 signal segments. Reached 99.96% (2-class), 90.73% (4-class), 70.09% (10-class). Overall stronger than Al-Sa’d et al. [3] for simpler tasks, but lags on higher-class classification vs. more recent solutions. |
Li et al. (2021) [72] | RF | Bispectrum Siamese Network + Contrastive Learning | DroneRF [8] | Converts UAV signals to bispectrum for unsupervised representation. Achieved 100% (2-class), 98.57% (4-class), and 92.31% (10-class). Gains up to 10% over SimCLR, SimSiam, and GAN baselines, though still below XGBoost for large class sets. |
Nemer et al. (2021) [86] | RF | Ensemble learning with hierarchical classification | DroneRF [8] | Multi-stage scheme yields 99% overall accuracy on ten-type classification. Outperforms earlier CNN approaches (e.g., [1]) for flight-mode tasks. |
Inani et al. (2023) [56] | RF | XGBoost + PSD features + 1DCNN | DroneRF [8] | Achieved 100% (2-class), 99.82% (4-class), 99.51% (10-class)—current state-of-the-art on DroneRF. Improves upon [80] and [3] across all class splits, leveraging PSD feature fusion with deep learning. |
@@
Al-Emadi et al. (2019) [2] | Aud. | CNN, RNN, CRNN | DroneAudio [2] | Released DroneAudio dataset with drone clips. Demonstrated CNN-based acoustic ID, achieving 96.38% (2-class) and 92.94% (3-class) accuracy. Provided a baseline for audio-based UAV detection. |
Katta et al. (2022) [61] | Aud. | CRNN, CNN, LSTM, CLSTM, Transformer | DroneAudio [2] | Benchmarked multiple neural architectures for UAV detection and ID on DroneAudio. LSTM outperformed CRNN/CNN with 98.93% accuracy (2-class) and 98.60% accuracy (3-class). |
Xiao et al. (2025) [148] | Aud. | CNN+RNN (temporal) | MMAUD (audio-only) [156] | Achieved 98.0% classification accuracy and 0.55 APE for 3D trajectory estimation with a Mamba-based audio pipeline (4-channel short melspectrograms). Demonstrated robust performance across day/night conditions. |
@@
Mendis et al. (2016) [81] | Radar | CNN (micro-Dopp.) | None public | Utilized deep belief networks for micro-Doppler SCF signatures, identifying UAV presence within radar beam width. One of the earliest radar-based UAV classification approaches, but lacks open data for direct replication. |
@@
Wisniewski et al. (2022) [139] | RGB | DenseNet201 CNN | Synthetic Drone [140] | Achieved 92.4% classification accuracy (distinguishing DJI Phantom, Mavic, and Inspire) with 88.8% precision, 88.6% recall, 88.7% F1-score. Demonstrated synthetic data viability for UAV type classification. |
@@
Deng et al. (2024) [32] | RGB, RA, L | Deep fusion | MMAUD [156] | Explored 4-class UAV classification in 3D pose estimation. Fused LiDAR, radar, and fisheye camera inputs via an EfficientNet-B7–based pipeline. Achieved 81.36% classification accuracy, improving keyframe-based multi-sensor approaches. |
Xiao et al. (2024) [147] | RGB, Aud. | AV-DTEC (self-supervised audio-visual fusion) | MMAUD [156] | Achieved 99.3% classification accuracy and 0.67m APE. Parallel state-space fusion boosted performance under varied illumination. Offers an alternative to full multi-sensor pipelines by only focusing on audio-visual synergy. |
@@ 2 UAV Datasets & Benchmarking
Foundational Efforts & Modality Expansion. Initial benchmarks such as FL-Drones [103], NPS-Drones [68], and USC Drone [88, 138] provided core RGB datasets for aerial detection and tracking. Beyond the visual domain, DroneRF [8] captures electromagnetic signatures for RF-based UAV classification, while DroneAudio [2] and Acoustic-UAV [21] focus on acoustic features. More recent efforts extend coverage to multi-modal data: Halmstad Drone [120, 121] fuses RGB, IR, and audio, and MMAUD [156] integrates radar and LiDAR. Community-driven competitions (e.g., the 2nd and 3rd Anti-UAV Workshops [162, 164] and the 4th Anti-UAV Challenge [89]) feature adversarial or cluttered scenarios that reflect real-world operational complexity, while UAVSwarm [133] and mDrone [165] emphasize large-scale swarm coordination or low-visibility tracking—both highlighting the need for robust multi-sensor fusion [102].
Complex Scenes & Modalities. Recent datasets target increasingly demanding environments and more granular annotations. For instance, Anti-UAV600 [171] and Anti-UAV410 [51] rely on infrared or thermal imaging to capture nighttime or harsh weather conditions, while Drone-vs-Bird [29] focuses on distinguishing UAVs from visually similar wildlife. Other repositories (e.g., Midgard [131] and Rosner [102]) explore indoor or GPS-denied settings, alongside multi-view or re-identification tasks [11, 59]. Simultaneously, annotations have expanded from bounding boxes to 3D coordinates and comprehensive trajectory information [156], aligning with the growing complexity of interception and multi-UAV maneuvers.
Trends & Open Challenges. Despite notable advances in dataset scope, several gaps persist. Collections like MMAUD [156] and Halmstad Drone [121] require precise sensor calibration and synchronization, yet standardized fusion protocols remain limited. While UAVSwarm [133] includes up to 23 drones, testing on larger fleets (50+) remains uncommon, constraining research on scaled cooperative tracking and interception. Moreover, stealth or jamming scenarios—emerging as key topics in anti-UAV forums [162, 164, 89]—are still sparsely represented in open-source datasets. High-resolution streams (e.g., Det-Fly [166] and AOT [10]) also challenge onboard systems in terms of storage and processing, prompting further research into compression and hardware acceleration [11, 10].
@@
Zhao et al. (2022) [165] | RA | CNN + LSTM | mDrone [165] | Exploits Doppler signals from propellers for real-time detection and 3D tracking at 10 Hz. Achieves 9 cm localization error, outperforming five classic radar baselines (e.g., 3D MUSIC) by 66% in accuracy. |
@@
Seidaliyeva et al. (2020) [110] | RGB | Background subtraction + MobileNetV2 | Drone-vs-Bird [29], BirdDetection [154] | Combines Drone-vs-Bird with a bird dataset to enhance UAV-vs-bird discrimination. Reached 0.56 F1 at IoU=0.5 (9 FPS). Performance is improved upon by works using more advanced YOLO variants (e.g., [31]). |
@@
Dadboud et al. (2021) [31] | RGB | YOLOv5x (PANet neck, mosaic augment) | Det-Fly [166], Drone-vs-Bird [29] | Combined Drone-vs-Bird + Det-Fly, achieving 0.98 mAP@0.5 with 0.96 recall. Improves upon [110] on Drone-vs-Bird, though different training splits limit exact comparison. |
@@
Sangam et al. (2023) [108] | RGB | CSPDarkNet-53 + VideoSwin | FL-Drones [103], NPS-Drones [68], AOT [10] | On FL-Drones: 0.89 precision, 0.85 recall, 0.87 F1, 0.84 AP@0.5. On NPS-Drones: 0.94 precision, 0.92 recall, 0.93 F1, 0.93 AP@0.5. On AOT: 0.85 precision, 0.76 recall, 0.80 F1, 0.82 AP@0.5. Comparable to Rebbapragada et al. [98] under similar setups. |
@@
Tu et al. (2024) [128] | RGB | YOLOv8 + FCM + FMM | DUT Anti-UAV[163] | AP@0.5=0.913, AP@0.5:0.95=0.612, below the best YOLO-based methods on this dataset (e.g., [53]) but highlights a lightweight, efficient design. |
@@
Cheng et al. (2024) [26] | RGB | YOLOv5s + LGFFM | DUT Anti-UAV [163] | mAP@0.5=0.943. Improves the recall rate of 2.96% and the precision of 1.41% compared with YOLOv5s. Slightly outperforms [128] on the same benchmark. Low-complexity approach with 4.4 ms/frame inference. |
@@
Bo et al. (2024) [16] | RGB | YOLOv7-tiny + InceptionNeXt + SPPFCSPC-SR + “Get-and-Send” | DUT Anti-UAV [163] | mAP@0.5=0.932 at 104 FPS. Balances fast inference and high accuracy, though slightly lower mAP@0.5 than [26] and [53]. |
@@
Elsayed et al. (2024) [38] | RGB | YOLOv6 + LERFNet + large-kernel LAM | DUT Anti-UAV [163] | mAP@0.5=0.936, with 0.953 precision. Surpasses [16] in accuracy, albeit at presumably higher resource use than the YOLOv7-tiny approach. |
@@
Huang et al. (2024) [53] | RGB | YOLOv8 + ghost conv + EMA + DCNv2 | Det-Fly [166], DUT Anti-UAV [163] | On DUT Anti-UAV, yields 0.971 mAP@0.5, highest among YOLO-based reports here ([128, 26, 16, 38]). On Det-Fly, obtains 0.934 mAP@0.5. |
@@
Sun et al. (2024) [118] | IR | YOLOv8 w/ triple-input, BiFormer attention, small-object layer | Hui [55], Anti-UAV [58], Anti-UAV410 [51] | On Hui dataset: 0.943 mAP@0.5 at 72 FPS. Generalizes well to Anti-UAV ( mAP@0.5) and Anti-UAV410 ( mAP@0.5). Demonstrates strong IR-based detection across multiple IR benchmarks. |
@@
Wisniewski et al. (2024) [141] | RGB | Faster R-CNN | MAV-VID [100], Anti-UAV [58], Drone-vs-Bird [29] | AP@0.5=0.970 on MAV-VID, AP@0.5=0.498 on Drone-vs-Bird, AP@0.5=0.678 on Anti-UAV. |
@@
Zhou et al. (2024) [169] | RGB | YOLOv8 + multi-scale spatial attention + HDC | Det-Fly [166] | Achieved 0.815 precision, 0.849 recall. Gains of +0.05 precision/+0.06 recall relative to YOLOv8 baseline. Lags behind [53] and [46] in absolute mAP on Det-Fly, but emphasizes robust small-object detection. |
@@
Rebbapragada et al. (2024) [98] | RGB | SWIN + DETR + OEN | FL-Drones [103], NPS-Drones [68], AOT [10] | Matches Sangam et al. [108] in precision/recall/AP@0.5 = 0.89/0.85/0.84 on FL-Drones and 0.94/0.92/0.93 on NPS-Drones with a transformer-based pipeline. On AOT dataset, obtains 0.85 precision, 0.76 recall, AP@0.5=0.82 comparable to [108]. |
@@
Hao et al. (2025) [46] | RGB | YOLOv8 w/ CSDE, CSP-PMSA, OmniKernel, DERS head | Det-Fly [166], MOT-Fly [28] | On MOT-Fly, hits mAP@0.5=0.949 with only 1.13M parameters. On Det-Fly, achieves 0.964 mAP@0.5—slightly exceeding [53] at 0.934, though using different training splits. |
@@
Svanström et al. (2021) [122] | RGB, IR, Aud. | YOLOv2 (visible) + MFCC (audio) + LSTM | Halmstad Drone [122] | Achieved F10.61–0.88 across different distance with thermal IR sensor, F10.69–0.86 with visible camera, F10.93 with audio detector . Fusion cut false alarms vs. single-sensor methods, illustrating multi-modal benefits. |
Wu et al. (2024) [143] | RGB, RA | Attention-based Spatiotemporal Fusion Net | None public | Achieved 93.0% precision, 83.7% recall, 88.7% mAP in occluded scenes at 20 FPS. Outperforms vision-only detectors by fusing radar signals, though dataset not publicly released. |

### v3

@@ Abstract
| Dataset | Task | Modality | Seq. | Frame | Annotation | Resolution | Rate (Hz) | T-Type | S-Type | Venue (Year) | ||||
| FL-Drones [rozantsev2016detecting] | D, T | RGB | 14 | 39K | Bbox | 640480 / 752480 | – | 1 | 2 | TPAMI (2016) | ||||
| NPS-Drones [li2016multi] | D, T | RGB | 50 | 70K | Bbox | 19201280 / 1280760 | – | 1 | – | IROS (2016) | ||||
| USC Drone [USC_Drone] | D, T | RGB | 60 | 20K | Bbox | 19201080 | 15 | 1 | – | APSIPA ASC (2017) | ||||
| DroneRF [allahham2019dronerf] | C, D | RF | 454 | 10M | Class | 14-bit (RF) | Less than 1K (RF) | 3 | – | Data in Brief (2019) | ||||
| DroneAudio [al2019droneaudio] | C, D | A | – | 1332 clips | Class | – | 16K (A) | 2 | – | IWCMC (2019) | ||||
| MAV-VID [mavvid2020] | D, T | RGB | – | 40K | Bbox | – | – | 1 | – | IEEE Access (2020) | ||||
| Hui [Hui1] | D, T | IR | 22 | 16K | Centroid | 256256 | 100 | 1 | 22 | CSD (2020) | ||||
| Acoustic-UAV [2021Casabiancaacostic] | C, D | A | – | 91K sec | Class | – | 22.05K (A) | 9 | 6 | Drones (2021) | ||||
| mDrone [zhao2021mdrone] | C, D | RA | 100 | 60K | 3D coords | – | 10 (RA) | 1 | – | ICRA (2021) | ||||
| Det-Fly [zheng2021air] | D | RGB | – | 13K | Bbox | 38402160 | 5 | 1 | 4 | RA-L (2021) | ||||
| MOT-Fly [chu2023experimental] | C, D | RGB | 16 | 11K | Bbox | 19201080 | – | 3 | – | ICUS (2021) | ||||
| Halmstad Drone [halmstad_drone] | D, T | RGB, IR, A | 650 | 203K | Bbox |
4 | 3 | ICPR (2021) | ||||
| Drone-vs-Bird [Coluccia2021DroneBird] | C, D | RGB | 77 | 105K | Bbox | 720576 – 38402160 | – | 8 | – | AVSS (2021) | ||||
| Anti-UAV [anti_uav] | D, T | RGB, IR | 318 | 297K | Bbox |
25 | – | 7 | TMM (2021) | ||||
| UAVSwarm [wang2022uavswarm] | D, T | RGB | 72 | 12K | Bbox | 446276 – 19191079 | – | 19 | 13 | Remote Sens. (2022) | ||||
| Synthetic Drone [Wisniewski2022dataset] | C | RGB | – | 2K | Class | 256256 | – | 4 | 10 | CORD (2022) | ||||
| DUT Anti-UAV [dut_anti_uav] | D, T | RGB | 20 | 35K | Bbox | 160240 – 37445616 | – | 1 | 7 | T-ITS (2022) | ||||
| AOT [aot2023] | D, T | IR | 4943 | 5.9M+ | Bbox | 24482048 | 10 | – | – | AWS (2023) | ||||
| Anti-UAV600 [anti_uav_600] | D, T | IR | 600 | 723K | Bbox | 640512 | 25 | 1 | – | Preprint (2023) | ||||
| Anti-UAV410 [Huang2023AntiUAV410] | D, T | IR | 410 | 438K | Bbox | 640512 | 25 | 1 | 7 | TPAMI (2024) | ||||
| MMAUD [yuan2024mmaud] | C, D, T | RGB, A, RA, L | 50 | 45K | Trajectory | 2560960 (RGB) | 30 (RGB), 41.8 (A), 15 (RA), 10 (L) | 6 | – | ICRA (2024) | ||||
| RGBT-Tiny [ying2025visible] | D, T | RGB, IR | 115 | 93K | Bbox | 640512 | 15 | 7 | 8 | TPAMI (2025) |
@@ 1 Introduction
| Al-Sa’d et al. (2019) [AlSaad2019] | RF | 3 DNNs (RF fingerprint) | DroneRF [allahham2019dronerf] | Introduced DroneRF (open-source, multiple flight modes). Developed three DNN classifiers for UAV presence, type, and flight mode. Achieved 99.7% (2-class), 84.5% (4-class), 46.8% (10-class). Forms an early baseline on DroneRF. |
| Al-Emadi et al. (2020) [AlEmadi2020cnn] | RF | CNN | DroneRF [allahham2019dronerf] | Achieved 99.8% (2-class), 85.8% (4-class), and 59.2% (10-class). Comparable to Al-Sa’d et al. [AlSaad2019] on 2-class tasks, showing CNN’s viability for RF-based identification. |
| Medaiyese et al. (2021) [Medaiyese2021] | RF | XGBoost pipeline | DroneRF [allahham2019dronerf] | Used 227 signal segments. Reached 99.96% (2-class), 90.73% (4-class), 70.09% (10-class). Overall stronger than Al-Sa’d et al. [AlSaad2019] for simpler tasks, but lags on higher-class classification vs. more recent solutions. |
| Li et al. (2021) [li2021bissiam] | RF | Bispectrum Siamese Network + Contrastive Learning | DroneRF [allahham2019dronerf] | Converts UAV signals to bispectrum for unsupervised representation. Achieved 100% (2-class), 98.57% (4-class), and 92.31% (10-class). Gains up to 10% over SimCLR, SimSiam, and GAN baselines, though still below XGBoost for large class sets. |
| Nemer et al. (2021) [nemer2021] | RF | Ensemble learning with hierarchical classification | DroneRF [allahham2019dronerf] | Multi-stage scheme yields 99% overall accuracy on ten-type classification. Outperforms earlier CNN approaches (e.g., [AlEmadi2020cnn]) for flight-mode tasks. |
| Inani et al. (2023) [Inani2023] | RF | XGBoost + PSD features + 1DCNN | DroneRF [allahham2019dronerf] | Achieved 100% (2-class), 99.82% (4-class), 99.51% (10-class)—current state-of-the-art on DroneRF. Improves upon [Medaiyese2021] and [AlSaad2019] across all class splits, leveraging PSD feature fusion with deep learning. |
@@
| Al-Emadi et al. (2019) [al2019droneaudio] | Aud. | CNN, RNN, CRNN | DroneAudio [al2019droneaudio] | Released DroneAudio dataset with drone clips. Demonstrated CNN-based acoustic ID, achieving 96.38% (2-class) and 92.94% (3-class) accuracy. Provided a baseline for audio-based UAV detection. |
| Katta et al. (2022) [Katta2022] | Aud. | CRNN, CNN, LSTM, CLSTM, Transformer | DroneAudio [al2019droneaudio] | Benchmarked multiple neural architectures for UAV detection and ID on DroneAudio. LSTM outperformed CRNN/CNN with 98.93% accuracy (2-class) and 98.60% accuracy (3-class). |
| Xiao et al. (2025) [xiao2025tame] | Aud. | CNN+RNN (temporal) | MMAUD (audio-only) [yuan2024mmaud] | Achieved 98.0% classification accuracy and 0.55 APE for 3D trajectory estimation with a Mamba-based audio pipeline (4-channel short melspectrograms). Demonstrated robust performance across day/night conditions. |
@@
| Mendis et al. (2016) [mendis2016deep] | Radar | CNN (micro-Dopp.) | None public | Utilized deep belief networks for micro-Doppler SCF signatures, identifying UAV presence within radar beam width. One of the earliest radar-based UAV classification approaches, but lacks open data for direct replication. |
@@
| Wisniewski et al. (2022) [Wisniewski2022] | RGB | DenseNet201 CNN | Synthetic Drone [Wisniewski2022dataset] | Achieved 92.4% classification accuracy (distinguishing DJI Phantom, Mavic, and Inspire) with 88.8% precision, 88.6% recall, 88.7% F1-score. Demonstrated synthetic data viability for UAV type classification. |
@@
| Deng et al. (2024) [deng2024multi] | RGB, RA, L | Deep fusion | MMAUD [yuan2024mmaud] | Explored 4-class UAV classification in 3D pose estimation. Fused LiDAR, radar, and fisheye camera inputs via an EfficientNet-B7–based pipeline. Achieved 81.36% classification accuracy, improving keyframe-based multi-sensor approaches. |
| Xiao et al. (2024) [xiao2024av] | RGB, Aud. | AV-DTEC (self-supervised audio-visual fusion) | MMAUD [yuan2024mmaud] | Achieved 99.3% classification accuracy and 0.67m APE. Parallel state-space fusion boosted performance under varied illumination. Offers an alternative to full multi-sensor pipelines by only focusing on audio-visual synergy. |
@@ 2 UAV Datasets & Benchmarking
Foundational Efforts & Modality Expansion. Initial benchmarks such as FL-Drones [rozantsev2016detecting], NPS-Drones [li2016multi], and USC Drone [USCDroneDatasetWeb, Wang2019USCDrone] provided core RGB datasets for aerial detection and tracking. Beyond the visual domain, DroneRF [allahham2019dronerf] captures electromagnetic signatures for RF-based UAV classification, while DroneAudio [al2019droneaudio] and Acoustic-UAV [2021Casabiancaacostic] focus on acoustic features. More recent efforts extend coverage to multi-modal data: Halmstad Drone [svanstrom2021dataset, HalmstadDatasetRepo] fuses RGB, IR, and audio, and MMAUD [yuan2024mmaud] integrates radar and LiDAR. Community-driven competitions (e.g., the 2nd and 3rd Anti-UAV Workshops [zhao20212nd, zhao20233rd] and the 4th Anti-UAV Challenge [AntiUAVChallenge2023]) feature adversarial or cluttered scenarios that reflect real-world operational complexity, while UAVSwarm [wang2022uavswarm] and mDrone [zhao2021mdrone] emphasize large-scale swarm coordination or low-visibility tracking—both highlighting the need for robust multi-sensor fusion [rosner2025multimodal].
Complex Scenes & Modalities. Recent datasets target increasingly demanding environments and more granular annotations. For instance, Anti-UAV600 [anti_uav_600] and Anti-UAV410 [Huang2023AntiUAV410] rely on infrared or thermal imaging to capture nighttime or harsh weather conditions, while Drone-vs-Bird [Coluccia2021DroneBird] focuses on distinguishing UAVs from visually similar wildlife. Other repositories (e.g., Midgard [viktor2020midgard] and Rosner [rosner2025multimodal]) explore indoor or GPS-denied settings, alongside multi-view or re-identification tasks [andle2022stanford, kalra2019dronesurf]. Simultaneously, annotations have expanded from bounding boxes to 3D coordinates and comprehensive trajectory information [yuan2024mmaud], aligning with the growing complexity of interception and multi-UAV maneuvers.
Trends & Open Challenges. Despite notable advances in dataset scope, several gaps persist. Collections like MMAUD [yuan2024mmaud] and Halmstad Drone [HalmstadDatasetRepo] require precise sensor calibration and synchronization, yet standardized fusion protocols remain limited. While UAVSwarm [wang2022uavswarm] includes up to 23 drones, testing on larger fleets (50+) remains uncommon, constraining research on scaled cooperative tracking and interception. Moreover, stealth or jamming scenarios—emerging as key topics in anti-UAV forums [zhao20212nd, zhao20233rd, AntiUAVChallenge2023]—are still sparsely represented in open-source datasets. High-resolution streams (e.g., Det-Fly [zheng2021air] and AOT [aot2023]) also challenge onboard systems in terms of storage and processing, prompting further research into compression and hardware acceleration [andle2022stanford, aot2023].
@@
| Zhao et al. (2022) [zhao2021mdrone] | RA | CNN + LSTM | mDrone [zhao2021mdrone] | Exploits Doppler signals from propellers for real-time detection and 3D tracking at 10 Hz. Achieves 9 cm localization error, outperforming five classic radar baselines (e.g., 3D MUSIC) by 66% in accuracy. |
@@
| Seidaliyeva et al. (2020) [Seidaliyeva2020] | RGB | Background subtraction + MobileNetV2 | Drone-vs-Bird [Coluccia2021DroneBird], BirdDetection [yoshihashi2018] | Combines Drone-vs-Bird with a bird dataset to enhance UAV-vs-bird discrimination. Reached 0.56 F1 at IoU=0.5 (9 FPS). Performance is improved upon by works using more advanced YOLO variants (e.g., [Dadboud2021yolov5]). |
@@
| Dadboud et al. (2021) [Dadboud2021yolov5] | RGB | YOLOv5x (PANet neck, mosaic augment) | Det-Fly [zheng2021air], Drone-vs-Bird [Coluccia2021DroneBird] | Combined Drone-vs-Bird + Det-Fly, achieving 0.98 mAP@0.5 with 0.96 recall. Improves upon [Seidaliyeva2020] on Drone-vs-Bird, though different training splits limit exact comparison. |
@@
| Sangam et al. (2023) [Sangam2023TransVisDrone] | RGB | CSPDarkNet-53 + VideoSwin | FL-Drones [rozantsev2016detecting], NPS-Drones [li2016multi], AOT [aot2023] | On FL-Drones: 0.89 precision, 0.85 recall, 0.87 F1, 0.84 AP@0.5. On NPS-Drones: 0.94 precision, 0.92 recall, 0.93 F1, 0.93 AP@0.5. On AOT: 0.85 precision, 0.76 recall, 0.80 F1, 0.82 AP@0.5. Comparable to Rebbapragada et al. [Rebbapragada2024C2FDrone] under similar setups. |
@@
| Tu et al. (2024) [tu2024] | RGB | YOLOv8 + FCM + FMM | DUT Anti-UAV[dut_anti_uav] | AP@0.5=0.913, AP@0.5:0.95=0.612, below the best YOLO-based methods on this dataset (e.g., [huang2024EDGSYOLOv8]) but highlights a lightweight, efficient design. |
@@
| Cheng et al. (2024) [Cheng2024LocalGlobal] | RGB | YOLOv5s + LGFFM | DUT Anti-UAV [dut_anti_uav] | mAP@0.5=0.943. Improves the recall rate of 2.96% and the precision of 1.41% compared with YOLOv5s. Slightly outperforms [tu2024] on the same benchmark. Low-complexity approach with 4.4 ms/frame inference. |
@@
| Bo et al. (2024) [bo2024yolov7gs] | RGB | YOLOv7-tiny + InceptionNeXt + SPPFCSPC-SR + “Get-and-Send” | DUT Anti-UAV [dut_anti_uav] | mAP@0.5=0.932 at 104 FPS. Balances fast inference and high accuracy, though slightly lower mAP@0.5 than [Cheng2024LocalGlobal] and [huang2024EDGSYOLOv8]. |
@@
| Elsayed et al. (2024) [Elsayed2024LERFNetAE] | RGB | YOLOv6 + LERFNet + large-kernel LAM | DUT Anti-UAV [dut_anti_uav] | mAP@0.5=0.936, with 0.953 precision. Surpasses [bo2024yolov7gs] in accuracy, albeit at presumably higher resource use than the YOLOv7-tiny approach. |
@@
| Huang et al. (2024) [huang2024EDGSYOLOv8] | RGB | YOLOv8 + ghost conv + EMA + DCNv2 | Det-Fly [zheng2021air], DUT Anti-UAV [dut_anti_uav] | On DUT Anti-UAV, yields 0.971 mAP@0.5, highest among YOLO-based reports here ([tu2024, Cheng2024LocalGlobal, bo2024yolov7gs, Elsayed2024LERFNetAE]). On Det-Fly, obtains 0.934 mAP@0.5. |
@@
| Sun et al. (2024) [Sun2024multiyolov8] | IR | YOLOv8 w/ triple-input, BiFormer attention, small-object layer | Hui [Hui1], Anti-UAV [anti_uav], Anti-UAV410 [Huang2023AntiUAV410] | On Hui dataset: 0.943 mAP@0.5 at 72 FPS. Generalizes well to Anti-UAV ( mAP@0.5) and Anti-UAV410 ( mAP@0.5). Demonstrates strong IR-based detection across multiple IR benchmarks. |
@@
| Wisniewski et al. (2024) [wisniewski2024fastrcnn] | RGB | Faster R-CNN | MAV-VID [mavvid2020], Anti-UAV [anti_uav], Drone-vs-Bird [Coluccia2021DroneBird] | AP@0.5=0.970 on MAV-VID, AP@0.5=0.498 on Drone-vs-Bird, AP@0.5=0.678 on Anti-UAV. |
@@
| Zhou et al. (2024) [zhou2024MssamHdc] | RGB | YOLOv8 + multi-scale spatial attention + HDC | Det-Fly [zheng2021air] | Achieved 0.815 precision, 0.849 recall. Gains of +0.05 precision/+0.06 recall relative to YOLOv8 baseline. Lags behind [huang2024EDGSYOLOv8] and [hao2025yolov8] in absolute mAP on Det-Fly, but emphasizes robust small-object detection. |
@@
| Rebbapragada et al. (2024) [Rebbapragada2024C2FDrone] | RGB | SWIN + DETR + OEN | FL-Drones [rozantsev2016detecting], NPS-Drones [li2016multi], AOT [aot2023] | Matches Sangam et al. [Sangam2023TransVisDrone] in precision/recall/AP@0.5 = 0.89/0.85/0.84 on FL-Drones and 0.94/0.92/0.93 on NPS-Drones with a transformer-based pipeline. On AOT dataset, obtains 0.85 precision, 0.76 recall, AP@0.5=0.82 comparable to [Sangam2023TransVisDrone]. |
@@
| Hao et al. (2025) [hao2025yolov8] | RGB | YOLOv8 w/ CSDE, CSP-PMSA, OmniKernel, DERS head | Det-Fly [zheng2021air], MOT-Fly [chu2023experimental] | On MOT-Fly, hits mAP@0.5=0.949 with only 1.13M parameters. On Det-Fly, achieves 0.964 mAP@0.5—slightly exceeding [huang2024EDGSYOLOv8] at 0.934, though using different training splits. |
@@
| Svanström et al. (2021) [halmstad_drone] | RGB, IR, Aud. | YOLOv2 (visible) + MFCC (audio) + LSTM | Halmstad Drone [halmstad_drone] | Achieved F10.61–0.88 across different distance with thermal IR sensor, F10.69–0.86 with visible camera, F10.93 with audio detector . Fusion cut false alarms vs. single-sensor methods, illustrating multi-modal benefits. |
| Wu et al. (2024) [Wu2024multi] | RGB, RA | Attention-based Spatiotemporal Fusion Net | None public | Achieved 93.0% precision, 83.7% recall, 88.7% mAP in occluded scenes at 20 FPS. Outperforms vision-only detectors by fusing radar signals, though dataset not publicly released. |